# Auditing Black-Box LLM APIs with a Rank-Based Uniformity Test

### v2

@@ 3 Problem Formulation
where ${\pi}_{\text{alt}}$ is an arbitrary alternative model (e.g., quantized or adversarially finetuned version of ${\pi}_{\text{ref}}$) and $q:\mathcal{X}\!\to\![0,1]$ is an unknown routing function that may depend on the prompt and even on past interaction history (e.g., routing based on prompt caching (Gu et al., 2025). When $q(x)$ is small but positive on a non-negligible fraction of typical prompts, the mixture in (3) significantly complicates discrimination; nevertheless, a sound detection procedure must retain statistical power under such adversarial conditions.
@@ 4.1 Rank-Based Uniformity Test (RUT)
Since $f(y,x)$ takes values in a discrete set, $F_{{\pi}_{\text{ref}}}$ is a step function. To ensure the rank statistic is continuously distributed under the null, we apply a randomized quantile residual (Dunn and Smyth, 1996) to extend the probability integral transform (David and Johnson, 1948) to discrete distributions. Specifically, we define the rank statistic as
$r_{\text{tgt}}:=F_{{\pi}_{\text{ref}}}(s_{\text{tgt}}^{-})+U\cdot\mathbb{P}\left(f(y,x)=s_{\text{tgt}}\right),\quad U\sim\text{Uniform}[0,1],$
where $F_{{\pi}_{\text{ref}}}(s_{\text{tgt}}^{-}):=\mathbb{P}\left(f(y,x)<s_{\text{tgt}}\right)$ is the left-limit of the CDF at $s_{\text{tgt}}$, and $\mathbb{P}(f(y,x)=s_{\text{tgt}})$ is the probability mass at $s_{\text{tgt}}$. Under the null hypothesis ${\pi}_{\text{tgt}}={\pi}_{\text{ref}}$, this rank statistic $r_{\text{tgt}}\in[0,1]$ is uniformly distributed under the null hypothesis.
Conversely, suppose that $r_{tgt}\sim\text{Uniform}[0,1]$ under the randomized quantile residual construction. Since the CDF $F_{{\pi}_{\text{ref}}}(\cdot\mid x)$ is stepwise and non-decreasing, a uniformly distributed $r_{tgt}$ implies that the score $s_{tgt}$ follows the same discrete distribution as $s_{\text{ref}}$. By injectivity of $f$, this further implies that $y_{\text{tgt}}\sim{\pi}_{\text{ref}}(\cdot\mid x;\varphi)$, and hence ${\pi}_{\text{tgt}}={\pi}_{\text{ref}}$.
Thus, with an injective score function $f$, testing the uniformity of $F_{\text{ref}}(f(y_{\text{tgt}},x)\mid x)$ offers a valid signal for distinguishing $\pi_{\text{tgt}}$ from $\pi_{\text{ref}}$.
Empirical approximation of $F_{\text{ref}}$.  In practice, it is intractable to build the true CDF $F_{\text{ref}}(\cdot\mid x)$. Instead, we approximate it using an empirical CDF from $m$ reference samples for each prompt.
@@
We then define the randomized rank statistic $r_{i}\in[0,1]$ as
@@
where $U_{i}\sim\text{Uniform}[0,1]$ is an independent random variable to break ties uniformly, and ensure $r_{i}$ is an unbiased estimator of $F_{\text{ref}}(s_{i}\mid x_{i})$.
@@
${\pi}_{\text{ref}}(\cdot\mid x;\varphi)\neq{\pi}_{\text{tgt}}(\cdot\mid x;\varphi)\quad\Rightarrow\quad\mathcal{L}(S_{{\pi}_{\text{ref}}})\neq\mathcal{L}(S_{{\pi}_{\text{tgt}}}).$
@@
Thus, we find the most discriminative score function through empirical experiments. In Section 4.2, we compare five candidate score functions—log-likelihood, token rank, log-rank, entropy, and the log-likelihood log-rank ratio (Su et al., 2023)—and find that log-rank is the most discriminative in practice for separating responses by ${\pi}_{\text{ref}}$ and ${\pi}_{\text{tgt}}$, and therefore adopt it in our uniformity test.
@@ 4.2 Score Function Selection
  • Log-rank: the logarithm of the token rank.
  • Entropy: predictive entropy of ${\pi}_{\text{ref}}(\cdot\mid x)$.
  • Log-likelihood log-rank ratio (LRR): the ratio between total log-likelihood and total log-rank of the tokens in response. (Su et al., 2023).
@@
To identify the most discriminative score function, we conduct a Monte Carlo evaluation consisting of $500$ independent trials. In each trial, we randomly select $10$ prompts from the WildChat (Zhao et al., 2024) dataset and sample $50$ completions per prompt from both ${\pi}_{\text{ref}}$ and ${\pi}_{\text{tgt}}$, using a fixed temperature of $0.5$ and a maximum length of $30$ tokens. For each candidate score function, we compute the average AUROC (Bradley, 1997) across the $10$ prompts for each trial, yielding a distribution of $500$ AUROC scores per function. Across different model comparisons, we find that log-rank consistently yields the most separable AUROC distribution from $0.5$, indicating the strongest discriminative power. Figure 2 shows an example comparing Gemma-2-9b-it with its 4-bit quantized variant. Based on these results, we select log-rank as the scoring function for our uniformity test. Complete AUROC results are provided in Appendix A.
@@ 5.4 Detecting SFT
We study the setting where the API provider fine-tunes a model on instruction-following data. Specifically, we fine-tune two base models—Llama-3.2-3B-Instruct and Mistral-7B-Instruct-v0.3—on benign and harmful instruction-following datasets. We use Alpaca (Taori et al., 2023) as the benign dataset and BeaverTails (Ji et al., 2023) for harmful question answering. Each model is fine-tuned on 500 samples from the respective dataset for 5 epochs using LoRA (Hu et al., 2021) with rank 64 and $\alpha=16$, a batch size of 32, and a learning rate of $1\times 10^{-4}$ on a single A100. For each fine-tuned checkpoint, we conduct the substitution experiment and compute the statistical power AUC of the detection methods.
@@ 5.5 Detecting Full Model Replacement
Figure 4: Statistical power AUC for detecting full model substitution using three methods. Each cell shows the AUC score between a reference and a target model. Diagonal values represent self-comparisons.

### v3

@@ 3 Problem Formulation
where ${\pi}_{\text{alt}}$ is an arbitrary alternative model (e.g., quantized or adversarially finetuned version of ${\pi}_{\text{ref}}$) and $q:\mathcal{X}\!\to\![0,1]$ is an unknown routing function that may depend on the prompt and even on past interaction history (e.g., routing based on prompt caching (Gu et al., 2025)). When $q(x)$ is small but positive on a non-negligible fraction of typical prompts, the mixture in (3) significantly complicates discrimination; nevertheless, a sound detection procedure must retain statistical power under such adversarial conditions.
@@ 4.1 Rank-Based Uniformity Test (RUT)
Since $f(y,x)$ takes values in a discrete set, $F_{{\pi}_{\text{ref}}}$ is a step function. To ensure the rank statistic is continuously distributed under the null hypothesis, we apply a randomized quantile residual (Dunn and Smyth, 1996) to extend the probability integral transform (David and Johnson, 1948) to discrete distributions. Specifically, we define the rank statistic as
$r_{\text{tgt}}:=F_{{\pi}_{\text{ref}}}(s_{\text{tgt}}^{-})+U\cdot\mathbb{P}\left(f(y,x)=s_{\text{tgt}}\right),\quad U\sim\text{Uniform}[0,1],$ (4)
where $F_{{\pi}_{\text{ref}}}(s_{\text{tgt}}^{-}):=\mathbb{P}\left(f(y,x)<s_{\text{tgt}}\right)$ is the left-limit of the CDF at $s_{\text{tgt}}$, and $\mathbb{P}(f(y,x)=s_{\text{tgt}})$ is the probability mass at $s_{\text{tgt}}$. Under the null hypothesis ${\pi}_{\text{tgt}}={\pi}_{\text{ref}}$, this rank statistic $r_{\text{tgt}}\in[0,1]$ is uniformly distributed.
Conversely, suppose that $r_{\text{tgt}}\sim\text{Uniform}[0,1]$ under the randomized quantile residual construction. Since the CDF $F_{{\pi}_{\text{ref}}}(\cdot\mid x)$ is stepwise and non-decreasing, a uniformly distributed $r_{\text{tgt}}$ implies that the score $s_{\text{tgt}}$ follows the same discrete distribution as $s_{\text{ref}}$. By injectivity of $f$, this further implies that $y_{\text{tgt}}\sim{\pi}_{\text{ref}}(\cdot\mid x;\varphi)$, and hence ${\pi}_{\text{tgt}}={\pi}_{\text{ref}}$.
Thus, with an injective score function $f$, testing the uniformity of $r_{\text{tgt}}$ as defined in (4) offers a valid signal for distinguishing $\pi_{\text{tgt}}$ from $\pi_{\text{ref}}$.
Empirical approximation of $F_{\pi_{\text{ref}}}$.  In practice, it is intractable to build the true CDF $F_{\pi_{\text{ref}}}(\cdot\mid x)$. Instead, we approximate it using an empirical CDF from $m$ reference samples for each prompt.
@@
We then define the randomized rank statistics $r_{i}\in[0,1]$ as
@@
where $U_{i}\sim\text{Uniform}[0,1]$ is an independent random variable to break ties uniformly, and ensure $r_{i}$ is an unbiased estimator of $r_{\text{tgt}}$ given the prompt $x_{i}$.
@@
${\pi}_{\text{ref}}(\cdot\mid x;\varphi)\neq{\pi}_{\text{tgt}}(\cdot\mid x;\varphi)\quad\Rightarrow\quad P_{S_{{\pi}_{\text{ref}}}}\neq P_{S_{{\pi}_{\text{tgt}}}}.$
@@
Thus, we aim to find the most discriminative score function among several promising candidates through empirical experiments. In Section 4.2, we compare five candidate score functions—log-likelihood, token rank, log-rank, entropy, and the log-likelihood log-rank ratio (Su et al., 2023)—and find that log-rank is the most discriminative in practice for separating responses by ${\pi}_{\text{ref}}$ and ${\pi}_{\text{tgt}}$, and therefore adopt it in our uniformity test.
@@ 4.2 Score Function Selection
  • Log-rank: the average of the logarithm of the token rank.
  • Entropy: predictive entropy for $y$ under ${\pi}_{\text{ref}}(x)$.
  • Log-likelihood log-rank ratio (LRR): the ratio between log-likelihood and log-rank. (Su et al., 2023).
@@
To identify the most discriminative score function, we conduct a Monte Carlo evaluation consisting of $500$ independent trials. In each trial, we randomly select $10$ prompts from the WildChat (Zhao et al., 2024) dataset and sample $50$ completions per prompt from both ${\pi}_{\text{ref}}$ and ${\pi}_{\text{tgt}}$, using a fixed temperature of $0.5$ and a maximum length of $30$ tokens. For each candidate score function, we compute the average AUROC (Bradley, 1997) across the $10$ prompts for each trial, yielding a distribution of $500$ AUROC scores per function. The full algorithm to calculate per score function average AUROC is included in Appendix A.1. Across different model comparisons, we find that log-rank consistently yields the most separable AUROC distribution from $0.5$, indicating the strongest discriminative power. Figure 2 shows an example comparing Gemma-2-9b-it with its 4-bit quantized variant. Based on these results, we select log-rank as the scoring function for our uniformity test. Complete AUROC results are provided in Appendix A.2.
@@ 5.4 Detecting SFT
We study the setting where the API provider fine-tunes a model on instruction-following data. Specifically, we fine-tune two base models—Llama-3.2-3B-Instruct and Mistral-7B-Instruct-v0.3—on benign and harmful instruction-following datasets. We use Alpaca (Taori et al., 2023) as the benign dataset and BeaverTails (Ji et al., 2023) for harmful question answering. Each model is fine-tuned on 500 samples from the respective dataset for 5 epochs using LoRA (Hu et al., 2021) with rank 64 and $\alpha=16$, a batch size of 32, and a learning rate of $1\times 10^{-4}$ on a single A100. For each checkpoint, we compute the statistical power AUC of the detection methods.
@@ 5.5 Detecting Full Model Replacement
Figure 4: Statistical power AUC for detecting full model replacement. Each cell shows the AUC score between a reference and a target model. Diagonal values represent self-comparisons.
@@ Appendix A AUROC
A.1 AUROC Algorithm
Input: Prompt set $\mathcal{D}$; models ${\pi}_{\text{ref}}$, ${\pi}_{\text{tgt}}$; decoding parameters $\varphi=(\tau,L)$, where $\tau$ is temperature and $L$ is the maximum generation length; number of prompts $n$; completions per prompt per model $m$; score functions $\{\delta_{1},\dots,\delta_{K}\}$.
Output: Mean AUROC per score function, denoted $\mu_{\mathrm{AUROC}}(\delta)$.
1
2Draw $\{x_{1},\dots,x_{n}\}\sim\text{Uniform}(\mathcal{D})$;
3
4for $i\in\{1,\dots,n\}$ do
5       $\{y_{\text{ref}}^{(j)}\}_{j=1}^{m}\sim{\pi}_{\text{ref}}(\cdot\mid x_{i};\varphi)$;
6       $\{y_{\text{tgt}}^{(j)}\}_{j=1}^{m}\sim{\pi}_{\text{tgt}}(\cdot\mid x_{i};\varphi)$;
7       $\mathcal{Y}_{i}\leftarrow\{y_{\text{ref}}^{(j)}\}\cup\{y_{\text{tgt}}^{(j)}\}$;
8       $L_{i}\leftarrow\{0\}^{m}\cup\{1\}^{m}$;
9
10      for $\delta\in\{\delta_{1},\dots,\delta_{K}\}$ do
11             $S_{i}\leftarrow\{\delta(y)\mid y\in\mathcal{Y}_{i}\}$;
12             Store $A_{i}^{\delta}\leftarrow\mathrm{AUROC}(S_{i},L_{i})$;
13
14
15
16for $\delta\in\{\delta_{1},\dots,\delta_{K}\}$ do
17       $\mu_{\mathrm{AUROC}}(\delta)\leftarrow\frac{1}{n}\sum_{i=1}^{n}A_{i}^{\delta}$;
18
Algorithm 1 Average AUROC for score function evaluation
Note. $\mathrm{AUROC}(S,L)$ denotes the standard binary AUROC (Bradley, 1997).
A.2 AUROC Score Distributions